### 05 Mar 2024 v1

@@ 5 The impact of numerical chaos
In general, numerical integrations are constrained by the inherent limitations of floating-point arithmetic precision, the temporal discretization of the solution, and the truncation error of the integration algorithm (see 52 for a discussion). The simulations presented thus far also suffer from these limitations (even though in tsunami the truncation errors are compensated by employing Richardson extrapolation, 53). Small integration errors accumulate over time and result in a deviation of our approximate solution from the mathematical one (i.e., the true mathematical solution that is guaranteed to exist by virtue of the Cauchy–Lipschitz theorem), indicating a lack of numerical convergence. Previous studies have highlighted that as long as numerical errors are below a certain threshold (impossible to determine a priori), an ensemble of approximate solutions can be statistically indistinguishable, even when individual solutions may not all be numerically converged [54].
Chaos influences the required accuracy and precision for simulations to achieve numerical convergence. In chaotic systems, small errors propagate and amplify much faster than in regular systems. Consequently, regular systems have less stringent numerical requirements for obtaining a converged solution [52]. The degree of chaos in a self-gravitating system depends on various factors. Generally, close encounters make a system more chaotic [55, 56, 57, 58], and systems with a larger number of particles have a shorter Lyapunov timescale [59, 60, 61, 62, 63, 64, 65, 66, 67, 68]. In the 3BP, chaos arises from the close encounters during democratic resonances.
@@
Since the term “numerical chaos” has only been introduced once before in the literature [69], it is essential to provide a clear definition to prevent confusion. Numerical chaos arises when numerical errors are amplified within a system characterized by both regular and chaotic regions in its phase space. For numerical chaos to occur, chaotic regions must exist, which diffuse throughout the phase space and infiltrate otherwise regular regions because of numerical errors. Put simply, numerical errors cannot induce chaos in a system that is entirely regular.
Consistent with the findings of Boekholt and Portegies Zwart [54], numerical chaos does not significantly affect the macroscopic outcomes of the interactions, with the difference in outcome statistics being less than 0.5% (Table 1). However, it is important to note that for this statement to hold true, the sample size must be sufficiently large.
@@
These results indicate that a small source of numerical noise has the same effect as a physical perturbation, which enables transport between neighbouring Hamiltonian flows [70]. Numerical chaos can therefore accelerate the rate at which an ensemble of orbits diffuse in the phase space, hiding the intrinsic regularity of a dynamical system.
@@
Columns from left to right: fraction of outcomes with escaper of mass $12.5$, $15$ or $17.5\,{{\rm M}_{\odot}}$, respectively. The bottom row shows the relative difference in percentage.
@@
00footnotetext: Columns from left to right: fraction of outcomes with escaper of mass $12.5$, $15$ or $17.5\,{{\rm M}_{\odot}}$, respectively. The last column indicates all regular interactions. The bottom row shows the relative difference in percentage.
@@ 6 Perturbations below the Planck scale
Figure 10: Space of the initial configurations, color-coded by ejected particle (left) and lifetime of the system (right). Each pixel corresponds to a realization on a $320\times 320$ grid. As in Figure 5, the axes denote binary phase and inclination, but since displaying the actual values in degrees would require too many digits, we just show the realization number on the grid. The size of the box is equivalent to a perturbation in the particles’s positions comparable to the Planck length. The simulations were run with brutus with a tolerance of $10^{-63}$ and a word-length of 356 bits.
@@
With compelling evidence demonstrating that chaos and regularity persist at any scale, including scales below the Planck length, any microscopic stochastic effect below the Planck length has the potential to propagate macroscopically [see also 71]. In this context, the regular behaviour of the 3BP provides a means to connect micro- and macroscopic physics.
@@ 7 Conclusions
One potential critique of our work could be that our results are only applicable to our specific three-body setup, which admittedly represents a small portion of the vast parameter space of the 3BP. To address this concern, we conducted additional numerical experiments involving hyperbolic binary-single scatterings with initial velocities at infinity and varied impact parameters. Remarkably, we observed the same regular patterns as depicted in Figure 3, indicating the robustness of our findings.
@@
Another crucial variation of parameters that we have not yet explored is the masses of the three bodies, or rather, their mass ratio. Intuitively, as the mass of the third body decreases, the motion of the two most massive bodies is expected to become increasingly regular. In the extreme regime of the restricted 3BP with $m_{3}\ll m_{1},m_{2}$, the binary moves unperturbed on a Keplerian orbit, while the motion of the test particle remains chaotic. Manwadkar et al. [27] showed that as the mass contrast between the most and least massive bodies increases, the regular islands depicted in Figure 3 begin to disintegrate. Exploring the transition between the equal mass and test particle limits, as well as the limitations of escape theories in this context, presents an interesting direction for future research.
@@
Our findings underscore that regular regions in the phase space pose challenges to the effectiveness of statistical escape theories as predictive tools for three-body interactions. Consequently, caution is advised when applying such tools to astrophysical problems [e.g. 42, 43, 72, 44]. Specifically, the eccentricity distribution of final binaries from regular trajectories diverges significantly from the expectations of statistical escape theories, which introduces a bias into gravitational wave population synthesis models reliant on such theories.
[…]
Acknowledgments
@@
AAT acknowledges support from JSPS KAKENHI Grant Number 21K13914 and from the European Union’s Horizon 2020 and Horizon Europe research and innovation programs under the Marie Skłodowska-Curie grant agreements no. 847523 and 101103134. All the simulations were performed on the awamori computer cluster at The University of Tokyo. TB is supported by an appointment to the NASA Postdoctoral Program at the NASA Ames Research Center, administered by Oak Ridge Associated Universities under contract with NASA. NWCL gratefully acknowledges the generous support of a Fondecyt General grant 1230082, as well as support from Núcleo Milenio NCN2023_002 (TITANs) and funding via the BASAL Centro de Excelencia en Astrofisica y Tecnologias Afines (CATA) grant PFB-06/2007. NWCL also thanks support from ANID BASAL project ACE210002 and ANID BASAL projects ACE210002 and FB210003. AAT, SPZ and TB also thank Anna Lisa Varri for organizing the “Chaotic rendezvous” meeting at the University of Edinburgh, where part of this work was completed.

### 08 Mar 2024 v2

@@ 5 The impact of numerical chaos
In general, numerical integrations are constrained by the inherent limitations of floating-point arithmetic precision, the temporal discretization of the solution, and the truncation error of the integration algorithm (see Portegies Zwart & Boekholt, 2018 for a discussion). The simulations presented thus far also suffer from these limitations (even though in tsunami the truncation errors are compensated by employing Richardson extrapolation, Press et al., 2007). Small integration errors accumulate over time and result in a deviation of our approximate solution from the mathematical one (i.e., the true mathematical solution that is guaranteed to exist by virtue of the Cauchy–Lipschitz theorem), indicating a lack of numerical convergence. Previous studies have highlighted that as long as numerical errors are below a certain threshold (impossible to determine a priori), an ensemble of approximate solutions can be statistically indistinguishable, even when individual solutions may not all be numerically converged (Boekholt & Portegies Zwart, 2015).
Chaos influences the required accuracy and precision for simulations to achieve numerical convergence. In chaotic systems, small errors propagate and amplify much faster than in regular systems. Consequently, regular systems have less stringent numerical requirements for obtaining a converged solution (Portegies Zwart & Boekholt, 2018). The degree of chaos in a self-gravitating system depends on various factors. Generally, close encounters make a system more chaotic (Suto, 1991; Portegies Zwart et al., 2023; Boekholt et al., 2023; Boekholt & Portegies Zwart, 2023), and systems with a larger number of particles have a shorter Lyapunov timescale (Heggie, 1988; Kandrup & Smith, 1991, 1992; Kandrup et al., 1992; Goodman et al., 1993; Kandrup et al., 1994; Hemsendorf & Merritt, 2002; Di Cintio & Casetti, 2019, 2020; Portegies Zwart et al., 2022). In the 3BP, chaos arises from the close encounters during democratic resonances.
@@
Since the term “numerical chaos” has only been introduced once before in the literature (Ablowitz et al., 1993), it is essential to provide a clear definition to prevent confusion. Numerical chaos arises when numerical errors are amplified within a system characterized by both regular and chaotic regions in its phase space. For numerical chaos to occur, chaotic regions must exist, which diffuse throughout the phase space and infiltrate otherwise regular regions because of numerical errors. Put simply, numerical errors cannot induce chaos in a system that is entirely regular.
Consistent with the findings of Boekholt & Portegies Zwart (2015), numerical chaos does not significantly affect the macroscopic outcomes of the interactions, with the difference in outcome statistics being less than 0.5% (Table 1). However, it is important to note that for this statement to hold true, the sample size must be sufficiently large.
@@
These results indicate that a small source of numerical noise has the same effect as a physical perturbation, which enables transport between neighbouring Hamiltonian flows (Kandrup & Novotny, 2004). Numerical chaos can therefore accelerate the rate at which an ensemble of orbits diffuse in the phase space, hiding the intrinsic regularity of a dynamical system.
@@
 [^1] Columns from left to right: fraction of outcomes with escaper of mass $12.5$, $15$ or $17.5\,{{\rm M}_{\odot}}$, respectively. The bottom row shows the relative difference in percentage.
@@
 [^2] Columns from left to right: fraction of outcomes with escaper of mass $12.5$, $15$ or $17.5\,{{\rm M}_{\odot}}$, respectively. The last column indicates all regular interactions. The bottom row shows the relative difference in percentage.
@@ 6 Perturbations below the Planck scale
Figure 10: Space of the initial configurations, color-coded by ejected particle (left) and lifetime of the system (right). Each pixel corresponds to a realization on a $320\times 320$ grid. As in Figure 5, the axes denote binary phase and inclination, but since displaying the actual values in degrees would require too many digits, we just show the realization number on the grid. The size of the box is equivalent to a perturbation in the particles’s positions comparable to the Planck length. The simulations were run with brutus with a tolerance of $10^{-63}$ and a word length of 356 bits.
@@
With compelling evidence demonstrating that chaos and regularity persist at any scale, including scales below the Planck length, any microscopic stochastic effect below the Planck length has the potential to propagate macroscopically (see also Boekholt et al., 2020). In this context, the regular behaviour of the 3BP provides a means to connect micro- and macroscopic physics.
7 Astrophysical implications
Statistical escape theories of the 3BP have recently gained traction in astrophysics due to their computational efficiency compared to direct three-body simulations. Notably, these theories have found various applications in predicting the properties of merging black holes resulting from dynamical interactions in dense stellar environments (e.g., Kritos et al., 2022; Mapelli et al., 2022; Ginat & Perets, 2023). Additionally, a recent study has utilized statistical escape theories to model the ejection of extra-tidal stars from globular clusters in the Galactic halo (Grondin et al., 2023).
Our findings raise concerns regarding the application of statistical escape theories to astrophysical scenarios. While all trajectories we calculated numerically are physically plausible and can occur in nature, statistical escape theories only predict those trajectories that exhibit sufficient chaotic behavior to mix the phase space. In our setup, this results in the exclusion of approximately 32% of the initial phase space. As demonstrated in Section 3, this exclusion biases the distribution of outcome parameters (such as binary semimajor axis and eccentricity, and single escape velocity), as well as the escape probability of the single, based on its mass.
In Figure 11, we illustrate how these biases impact the estimation of the gravitational wave coalescence time $t_{\rm gw}$ for the final binaries in our simulations, assuming they are black holes (Peters, 1964). Despite constituting only about a third of the total, binaries formed from regular interactions exhibit significantly shorter coalescence times compared to those formed from chaotic interactions, because of their higher eccentricity. Moreover, binaries resulting from regular interactions with short ($t_{\rm gw}<10^{7}\rm\,yr$) coalescence times are roughly twice as numerous as those from chaotic interactions. Our analysis therefore suggests that statistical escape theories overestimate the coalescence time, and thus underestimate the merger efficiency of binaries formed from three-body interactions. Another likely consequence of the bias in eccentricity is the underestimation of the fraction of eccentric mergers that may be detectable with gravitational wave interferometers (Abbott et al., 2019; Romero-Shaw et al., 2021; Gayathri et al., 2022; The LIGO Scientific Collaboration et al., 2023).
Figure 11: Top panel: ratio of the cumulative number of binaries formed from regular interactions to that of binaries formed from chaotic interactions, as a function of gravitational wave coalescence time $t_{\rm gw}$. Bottom panel: coalescence time distribution of final binaries for the regular ($f_{\rm col}\leq 2/3$, olive) and chaotic ($f_{\rm col}>2/3$, purple) interactions. Both distributions are normalized to the total.
We emphasize that the fraction of regular trajectories, as well as their outcome properties, should vary depending on the specific initial setup of the 3BP. Nevertheless, we anticipate the coexistence of regular and chaotic regions to be a natural feature of the 3BP, as supported by findings in other studies with different setups (e.g. Manwadkar et al., 2021; Parischewsky et al., 2023; Manwadkar et al., 2023). For example, Manwadkar et al. (2023) investigated the ’chaotic absorptivity’ (the probability for an interaction to enter chaotic motion) for hyperbolic binary-single encounters with varying initial binary angular momentum and energy. Their figure 5 illustrates the persistent separation between chaotic and regular motion in the energy-angular momentum space, with the fraction of chaotic phase space decreasing for larger initial binary separations. Thus, it is important to evaluate the biases introduced by the use of statistical escape theories to model astrophysical scenarios on a case-by-case basis. A comprehensive assessment of these biases is beyond the scope of this paper.
@@ 8 Conclusions
One potential critique of our work could be that our results are only applicable to our specific three-body setup, which admittedly represents a small portion of the vast parameter space of the 3BP. To address this concern, we conducted additional numerical experiments involving hyperbolic binary-single scatterings with an initial velocity at infinity and impact parameter. Remarkably, we observed the same regular patterns as depicted in Figure 3, indicating the robustness of our findings. Moreover, our conclusions are supported by the results of Manwadkar et al. (2023), who found that chaotic and regular motion occupy distinct regions in the energy-angular momentum space of hyperbolic binary-single encounters, even when statistically averaged over the binary phase.
@@
Another crucial variation of parameters that we have not yet explored is the masses of the three bodies, or rather, their mass ratio. Intuitively, as the mass of the third body decreases, the motion of the two most massive bodies is expected to become increasingly regular. In the extreme regime of the restricted 3BP with $m_{3}\ll m_{1},m_{2}$, the binary moves unperturbed on a Keplerian orbit, while the motion of the test particle remains chaotic. Manwadkar et al. (2020) showed that as the mass contrast between the most and least massive bodies increases, the regular islands depicted in Figure 3 begin to disintegrate. Exploring the transition between the equal mass and test particle limits, as well as the limitations of escape theories in this context, presents an interesting direction for future research.
@@
Our findings underscore that regular regions in the phase space pose challenges to the effectiveness of statistical escape theories as predictive tools for three-body interactions. Consequently, caution is advised when applying such tools to astrophysical problems (e.g. Kritos et al., 2022; Mapelli et al., 2022; Grondin et al., 2023; Ginat & Perets, 2023). Specifically, the eccentricity distribution of final binaries from regular trajectories diverges significantly from the expectations of statistical escape theories, which introduces a bias into gravitational wave population synthesis models reliant on such theories.
Acknowledgements.
@@
AAT acknowledges support from JSPS KAKENHI Grant Number 21K13914 and from the European Union’s Horizon 2020 and Horizon Europe research and innovation programs under the Marie Skłodowska-Curie grant agreements no. 847523 and 101103134. All the simulations were performed on the awamori computer cluster at The University of Tokyo. TCNB is supported by an appointment to the NASA Postdoctoral Program at the NASA Ames Research Center, administered by Oak Ridge Associated Universities under contract with NASA. NWCL gratefully acknowledges the generous support of a Fondecyt General grant 1230082, as well as support from Núcleo Milenio NCN2023_002 (TITANs) and funding via the BASAL Centro de Excelencia en Astrofisica y Tecnologias Afines (CATA) grant PFB-06/2007. NWCL also thanks support from ANID BASAL project ACE210002 and ANID BASAL projects ACE210002 and FB210003. AAT, SPZ and TCNB also thank Anna Lisa Varri for organizing the “Chaotic rendezvous” meeting at the University of Edinburgh, where part of this work was completed. AAT thanks Mario Pasquato for being an accessible source of statistical knowledge at any time of the day.